# Sparse Variational Student-t Processes

### v1

@@ Introduction
Gaussian Processes (GPs) are widely used in regression tasks due to their flexibility and interpretability. A GP is a collection of random variables, any finite number of which have a joint Gaussian distribution. It is defined by a mean function $m(\mathbf{x})$ and a covariance function $k(\mathbf{x},\mathbf{x}^{\prime})$, which specify the distribution of the GP at any point in its input space.
For a regression problem, we observe a set of input-output pairs $\{\mathbf{x}_{i},y_{i}\}_{i=1}^{n}$, where $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ and $\epsilon_{i}$ is the noise term. We assume that $f(\mathbf{x})$ follows a GP with mean function $m(\mathbf{x})$ and covariance function $k(\mathbf{x},\mathbf{x}^{\prime})$. Then, the predictive distribution of the function values at a new test point $\mathbf{x}^{*}$ is given by:
@@
To address this issue, sparse representation methods (Hensman, Matthews, and Ghahramani 2015) have been proposed, which use a small number of inducing points $\mathbf{Z}$ to approximate the full covariance matrix of the GP. Sparse Gaussian processes can be formulated using various techniques. One popular approach is the Sparse Variational Gaussian processes (SVGP), which use a variational distribution to approximate the true GP posterior. In SVGP, we introduce a set of $m$ inducing points $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$ and corresponding function values $\mathbf{u}=\{u_{i}\}_{i=1}^{m}$ at these points. Then, we introduce a variational distribution $q(\mathbf{u})$ to approximate the true posterior distribution $p(\mathbf{u}|\mathbf{y})$ of the inducing function values. The choice of the variational distribution $q(\mathbf{u})$ is often tractable, such as a Gaussian distribution.
Using the variational distribution $q(\mathbf{u})$, the predictive distribution at test points $\mathbf{x}^{*}$ is given by:
[…]
@@
where Γ()normal-$\Gamma(\cdot)$ is the gamma function, $\mathbf{y}\in\mathbb{R}^{n}$, and ||$\cdot$|| ⋅ | denotes the determinant of a matrix.
@@
[…]
@@
In this section, we propose the Sparse Student-t Process and develop a variational inference method for posterior sampling and prediction. We first define inducing points for TP and construct a variational lower bound for optimization and then present posterior sampling and prediction algorithms by gradient-based optimization.
@@
$p(\mathbf{u},\mathbf{f})=\mathcal{ST}\left(\nu,\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\end{bmatrix},\begin{bmatrix}K_{Z,Z^{\prime}}&K_{Z,X}\\ K_{X,Z}&K_{{X,X^{\prime}}}\end{bmatrix}\right)$ (6)
@@
[…]
@@
As Eq. (8) doesn’t restrict the distribution family of $q(\mathbf{u})$, theoretically any analytically parameterized distribution family can be used to parameterize the variational distribution. We propose to approximate $q(\mathbf{u})$ as a Student-t distribution, i.e. $q(\mathbf{u})=\mathcal{ST}(\widetilde{\nu},\mathbf{m},\mathbf{S})$ Although the true posterior distribution is intractable, we intuitively consider that the posterior is closer to a Student’s t-distribution because both the prior $p(\mathbf{u})$ and conditional distribution $p(\mathbf{f}|\mathbf{u})$ are Student’s t-distributions.
We rephrase the Eq. (8) and summarize the two terms of our variational lower bound $\mathcal{L}(q)$. The statistical interpretations of these terms are the expected likelihood function and KL regularization. Next, we will compute and optimize these two terms separately.
@@
$\mathbb{E}_{q(\mathbf{u})}[\log q(\mathbf{u})-\log p(\mathbf{u})]$ (12)
$=\underset{C\left(\nu,\tilde{\nu},\mathbf{S}\right)}{\underbrace{\begin{array}[]{c}\frac{1}{2}\log\frac{|K_{Z,Z^{\prime}}|}{|\mathbf{S}|}+\frac{M}{2}\log\frac{\nu-2}{\tilde{\nu}-2}+\log\Gamma\left(\frac{\tilde{\nu}+M}{2}\right)\\ -\log\Gamma\left(\frac{\tilde{\nu}}{2}\right)-\log\Gamma\left(\frac{\nu+M}{2}\right)+\log\Gamma\left(\frac{\nu}{2}\right)\\ \end{array}}}$
$-\frac{\tilde{\nu}+M}{2}\underset{\mathcal{L}_{1}}{\underbrace{\mathbb{E}_{q(\mathbf{u})}\left\{\log\left[1+\frac{1}{\tilde{\nu}-2}\left(\mathbf{u}-\mathbf{m}\right)^{\mathrm{T}}\mathbf{S}^{-1}\left(\mathbf{u}-\mathbf{m}\right)\right]\right\}}}$
@@
Eq. (12) requires the computation of two expectations, $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, which represent the entropy of $q(\mathbf{u})$ and the cross-entropy between $q(\mathbf{u})$ and $p(\mathbf{u})$, respectively  [^1] In this context, we omit some constant terms in the definition of entropy.. First, for $\mathcal{L}_{1}$, we can establish its relationship with the standard Student-t distribution using the following lemma:
@@
Next, we can use the following lemma to demonstrate that the entropy of the standard student-t distribution can be expressed in terms of the gamma function (and others 1999; Villa and Rubio 2018) with respect to $\nu$:
@@
where Ψ()normal-$\Psi(\cdot)$ is the digamma function, i.e. $\Psi\left(x\right)=\frac{d\log\varGamma\left(x\right)}{dx}$ for $x\in\mathbb{R}$.
@@
Although this approach introduces additional noise, it serves to accelerate convergence. Furthermore, this method is applicable to any likelihood function $p(y_{i}|f_{i})$.
@@
where $\beta=\mathbf{u}^{T}K_{Z,Z^{\prime}}^{-1}\mathbf{u}$, $\boldsymbol{k}_{*}$ denotes the kernel vector between inducing points $Z$ and the new input points $X_{*}$, $k_{**}$ is the kernel between the new input points. We can finally obtain the predicted distribution $\mathbf{f}^{*}\sim\mathcal{ST}(\nu+M,\mu_{*},\Sigma_{*})$, $\mathbf{y}^{*}=\mathbf{f}^{*}+\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}=\left\{\epsilon_{i}\right\}_{i=1}^{N}$
Specifically, we have discovered an interesting fact that when we set $q(\mathbf{u})=\mathcal{ST}(\nu+n,\mathbf{m},\mathbf{S})$ , we can analytically marginalize $\mathbf{u}$, resulting in (we proof it in the appendix):
[…]
@@
When discussing the differences between SVTP and SVGP in handling outliers in regression analysis, we can initially ignore the influence of the noise term $\epsilon_{i}$ and assume that $y_{i}=f_{i}$. Following (Hensman, Matthews, and Ghahramani 2015), the variational lower bound for SVGP can be expressed as:
@@
We evaluate the performance of our proposed methods, SVTP-UB and SVTP-MC, on a total of eight real-world datasets obtained from UCI and Kaggle. These datasets are used for comparison against two baseline models, namely Sparse Variational Gaussian Processes (SVGP) and the full Student-T Processes (TP).
Our experiments are designed to verify the following four propositions:
@@
Table 1: Comparison of Computational Complexity on regression experiments. In this context, $m$ refers to the number of inducing points. "--" stands for not having enough space and computational power to calculate its time.
@@
Table 2: Predictive Mean Squared Errors (MSE) and test Log Likelihoods (LL) of regression experiments
Dataset MSE LL
@@
Yacht $7.56\pm 0.23$ $\mathbf{1.44}\pm 0.02$ $1.55\pm 0.03$ $-485\pm 2.4$ $\mathbf{-114}\pm 0.5$ $-122\pm 0.7$
Energy $2.02\pm 0.06$ $\mathbf{0.74}\pm 0.04$ $0.84\pm 0.05$ $-156\pm 1.8$ $\mathbf{-63.3}\pm 1.2$ $-70.5\pm 1.3$
Boston $3.17\pm 0.21$ $\mathbf{1.82}\pm 0.10$ $2.01\pm 0.12$ $-171\pm 4.8$ $-\mathbf{126}\pm 2.6$ $-140\pm 2.9$
Elevator $6.54\pm 0.28$ $1.93\pm 0.15$ $\mathbf{1.84}\pm 0.11$ $-391\pm 19.2$ $-126.1\pm 10.2$ $-\mathbf{106}\pm 8.7$
Concrete $7.48\pm 0.24$ $5.71\pm 0.20$ $\mathbf{3.23}\pm 0.18$ $-478\pm 11.4$ $-435\pm 9.0$ $\mathbf{-276}\pm 7.4$
Protein $7.23\pm 0.26$ $5.64\pm 0.22$ $\mathbf{4.86}\pm 0.14$ $-425\pm 19.3$ $-295\pm 15.5$ $\mathbf{-260}\pm 9.8$
Kin8nm $5.19\pm 0.52$ $4.35\pm 0.47$ $\mathbf{3.85}\pm 0.46$ $-400\pm 28.6$ $-385\pm 22.5$ $-\mathbf{225}\pm 19.4$
Taxi $0.76\pm 0.07$ $0.54\pm 0.05$ $\mathbf{0.41}\pm 0.04$ $-301\pm 8.9$ $-152\pm 6.4$ $-\mathbf{124}\pm 4.6$
Concrete__\__Outliers $12.20\pm 0.32$ $8.62\pm 0.25$ $\mathbf{6.40}\pm 0.22$ $-658\pm 20.8$ $-471.0\pm 14.2$ $-\mathbf{336.3}\pm 13.6$
Kin8nm__\__Outliers $14.66\pm 2.25$ $9.12\pm 1.72$ $\mathbf{8.86}\pm 0.65$ $-1169\pm 61.9$ $-622\pm 52.4$ $-\mathbf{501}\pm 31.4$

### v2

@@ Introduction
Gaussian Processes (GPs) are widely used in regression tasks due to their flexibility and interpretability. A GP is a collection of random variables, any finite number of which have a joint Gaussian distribution. For a regression problem, we observe a set of input-output pairs $\{\mathbf{x}_{i},y_{i}\}_{i=1}^{n}$, where $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ and $\epsilon_{i}$ is the noise term. We assume that $f(\mathbf{x})$ follows a GP with mean function $m(\mathbf{x})$ and covariance function $k(\mathbf{x},\mathbf{x}^{\prime})$. Then, the predictive distribution of the function values at a new test point $\mathbf{x}^{*}$ is given by:
@@
To address this issue, sparse representation methods (Hensman, Matthews, and Ghahramani 2015) have been proposed, which use a small number of inducing points $\mathbf{Z}$ to approximate the full covariance matrix of the GP. Sparse Gaussian processes can be formulated using various techniques. One popular approach is the Sparse Variational Gaussian processes (SVGP), which use a variational distribution to approximate the true GP posterior. In SVGP, we introduce a set of $m$ inducing points $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$ and corresponding function values $\mathbf{u}=\{u_{i}\}_{i=1}^{m}$ at these points. Then, we introduce a variational distribution $q(\mathbf{u})$ to approximate the true posterior distribution $p(\mathbf{u}|\mathbf{y})$ of the inducing function values. The choice of the variational distribution $q(\mathbf{u})$ is often tractable, such as a Gaussian distribution. Using the variational distribution $q(\mathbf{u})$, the predictive distribution at test points $\mathbf{x}^{*}$ is given by:
$p(f(\mathbf{x}^{*})|\mathbf{x}^{*},\mathbf{X},\mathbf{y})=\int p(f(\mathbf{x}^{*})|\mathbf{x}^{*},\mathbf{Z},\mathbf{u})q(\mathbf{u})d\mathbf{u},$ (2)
@@
where $\Gamma(\cdot)$ is the gamma function, $\mathbf{y}\in\mathbb{R}^{n}$, and $|\cdot|$ denotes the determinant of a matrix.
@@
Then, given $y_{1}$, the conditional distribution of $y_{2}$ can be expressed as $\mathcal{S}\mathcal{T}_{n_{2}}(\nu+n_{1},\phi_{2}^{*},\frac{\nu+\beta_{1}-2}{\nu+n_{1}-2}t^{*})$, where $\phi_{2}^{*}=K_{21}K_{11}^{-1}(y_{1}-\phi_{1})+\phi_{2},\beta_{1}=(y_{1}-\phi_{1})^{T}K_{11}^{-1}(y_{1}-\phi_{1})$ and $t^{*}=K_{22}-K_{21}K_{11}^{-1}K_{12}$. The mean and covariance matrix of the conditional distribution of $y_{2}$ given $y_{1}$ are $\mathbb{E}[y_{2}|y_{1}]=\phi_{2}^{*}$ and $\text{cov}[y_{2}|y_{1}]=\frac{\nu+\beta_{1}-2}{\nu+n_{1}-2}t^{*}$, respectively.
@@
[…]
@@
$\log p(\mathbf{y})$ $=\log\int p(\mathbf{y},\mathbf{f},\mathbf{u})d\mathbf{f}d\mathbf{u}\geq\mathcal{L}(q)$ (8)
$=\int p(\mathbf{f}|\mathbf{u})q(\mathbf{u})\log\left(\frac{p(\mathbf{y}|\mathbf{f})p(\mathbf{f}|\mathbf{u})p(\mathbf{u})}{p(\mathbf{f}|\mathbf{u})q(\mathbf{u})}\right)d\mathbf{f}d\mathbf{u}$
@@
As Eq. (8) doesn’t restrict the distribution family of $q(\mathbf{u})$, theoretically any analytically parameterized distribution family can be used to parameterize the variational distribution. We propose to approximate $q(\mathbf{u})$ as a Student-t distribution, i.e. $q(\mathbf{u})=\mathcal{ST}(\widetilde{\nu},\mathbf{m},\mathbf{S})$ . Although the true posterior distribution is intractable, we intuitively consider that the posterior is closer to a Student’s t-distribution because both the prior $p(\mathbf{u})$ and conditional distribution $p(\mathbf{f}|\mathbf{u})$ are Student’s t-distributions.
We rephrase the Eq. (8) and summarize the two terms of our variational lower bound $\mathcal{L}(q)$. The statistical interpretations of these terms are the expected likelihood function and KL regularization.
@@
$\mathbb{E}_{q(\mathbf{u})}[\log q(\mathbf{u})-\log p(\mathbf{u})]=\frac{1}{2}\log\frac{|K_{Z,Z^{\prime}}|}{|\mathbf{S}|}$ (12)
$+\frac{M}{2}\log\frac{\nu-2}{\tilde{\nu}-2}+\log\Gamma\left(\frac{\tilde{\nu}+M}{2}\right)-\log\Gamma\left(\frac{\tilde{\nu}}{2}\right)$
$-\log\Gamma\left(\frac{\nu+M}{2}\right)+\log\Gamma\left(\frac{\nu}{2}\right)-\frac{\tilde{\nu}+M}{2}\cdot$
$\underset{\mathcal{L}_{1}}{\underbrace{\mathbb{E}_{q(\mathbf{u})}\left\{\log\left[1+\frac{1}{\tilde{\nu}-2}\left(\mathbf{u}-\mathbf{m}\right)^{\mathrm{T}}\mathbf{S}^{-1}\left(\mathbf{u}-\mathbf{m}\right)\right]\right\}}}$
@@
Eq. (12) requires the computation of two expectations, $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, which represent the entropy of $q(\mathbf{u})$ and the cross-entropy between $q(\mathbf{u})$ and $p(\mathbf{u})$, respectively . First, for $\mathcal{L}_{1}$, we can establish its relationship with the standard Student-t distribution using the following lemma:
@@
Next, we can use the following lemma to demonstrate that the entropy of the standard student-t distribution can be expressed in terms of the gamma function (Zografos et al. 1999; Villa and Rubio 2018) with respect to $\nu$:
@@
where $\Psi(\cdot)$ is the digamma function, i.e. $\Psi\left(x\right)=\frac{d\log\varGamma\left(x\right)}{dx}$ for $x\in\mathbb{R}$.
@@
where $\beta=\mathbf{u}^{T}K_{Z,Z^{\prime}}^{-1}\mathbf{u}$, $\boldsymbol{k}_{*}$ denotes the kernel vector between inducing points $Z$ and the new input points $X_{*}$, $k_{**}$ is the kernel between the new input points. We can finally obtain the predicted distribution $\mathbf{f}^{*}\sim\mathcal{ST}(\nu+M,\mu_{*},\Sigma_{*})$, $\mathbf{y}^{*}=\mathbf{f}^{*}+\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}=\left\{\epsilon_{i}\right\}_{i=1}^{N}$ .
Specifically, we have discovered an interesting fact that when we set $q(\mathbf{u})=\mathcal{ST}(\nu+n,\mathbf{m},\mathbf{S})$ , we can analytically marginalize $\mathbf{u}$, resulting in:
$q\left(\mathbf{f}\right)=\int{p\left(\mathbf{f}|\mathbf{u}\right)q\left(\mathbf{u}\right)}d\mathbf{u}=\mathcal{ST}\left(\nu,\mu^{\prime},\Sigma^{\prime}\right)$ (22)
@@
When discussing the differences between SVTP and SVGP in handling outliers in regression analysis, following (Hensman, Matthews, and Ghahramani 2015), the variational lower bound for SVGP can be expressed as:
@@
We evaluate the performance of our proposed methods, SVTP-UB and SVTP-MC, on a total of eight real-world datasets obtained from UCI and Kaggle. These datasets are used for comparison against two baseline models, namely Sparse Variational Gaussian Processes (SVGP) and the full Student-T Processes (TP). Our experiments are designed to verify the following four propositions:
@@
Table 1: Comparison of Computational Complexity on regression experiments. In this context, $m$ refers to the number of inducing points. ”-” stands for not having enough space and computational power to calculate its time.
@@
Dataset MSE $L$
@@
Yacht $7.56\pm 0.23$ $\mathbf{1.44}\pm 0.02$ $1.55\pm 0.03$ $485\pm 2.4$ $\mathbf{-114}\pm 0.5$ $122\pm 0.7$
Energy $2.02\pm 0.06$ $\mathbf{0.74}\pm 0.04$ $0.84\pm 0.05$ $156\pm 1.8$ $\mathbf{-63.3}\pm 1.2$ $70.5\pm 1.3$
Boston $3.17\pm 0.21$ $\mathbf{1.82}\pm 0.10$ $2.01\pm 0.12$ $171\pm 4.8$ $\mathbf{126}\pm 2.6$ $140\pm 2.9$
Elevator $6.54\pm 0.28$ $1.93\pm 0.15$ $\mathbf{1.84}\pm 0.11$ $391\pm 19.2$ $126.1\pm 10.2$ $\mathbf{106}\pm 8.7$
Concrete $7.48\pm 0.24$ $5.71\pm 0.20$ $\mathbf{3.23}\pm 0.18$ $478\pm 11.4$ $435\pm 9.0$ $\mathbf{-276}\pm 7.4$
Protein $7.23\pm 0.26$ $5.64\pm 0.22$ $\mathbf{4.86}\pm 0.14$ $425\pm 19.3$ $295\pm 15.5$ $\mathbf{-260}\pm 9.8$
Kin8nm $5.19\pm 0.52$ $4.35\pm 0.47$ $\mathbf{3.85}\pm 0.46$ $400\pm 28.6$ $385\pm 22.5$ $\mathbf{225}\pm 19.4$
Taxi $0.76\pm 0.07$ $0.54\pm 0.05$ $\mathbf{0.41}\pm 0.04$ $301\pm 8.9$ $152\pm 6.4$ $\mathbf{124}\pm 4.6$
Concrete_\_Outliers $12.20\pm 0.32$ $8.62\pm 0.25$ $\mathbf{6.40}\pm 0.22$ $658\pm 20.8$ $471.0\pm 14.2$ $\mathbf{336.3}\pm 13.6$
Kin8nm_\_Outliers $14.66\pm 2.25$ $9.12\pm 1.72$ $\mathbf{8.86}\pm 0.65$ $1169\pm 61.9$ $622\pm 52.4$ $\mathbf{501}\pm 31.4$
Table 2: Predictive Mean Squared Errors (MSE) and test Log Likelihoods ($L$) of regression experiments